# Improving Sequential Market Coordination via Value-oriented Renewable Energy Forecasting

### v2

@@ IV-A The Impact of the Forecast on the DA and RT Solutions
In addition, we define an operator $\Pi_{\mathcal{J}}:h(\bm{x})\mapsto\tilde{h}(\bm{x})$, where $h(\bm{x})=\bm{A}\bm{x}+\bm{b}$ is a linear function with the parameter $\bm{A},\bm{b}$, and $\mathcal{J}$ is the row index subset of $\bm{A},\bm{b}$. The output $\tilde{h}(\bm{x})$ of the operator is also a linear function, where $\tilde{h}(\bm{x})=\bm{A}[\mathcal{J}]\bm{x}+\bm{b}[\mathcal{J}]$ and $\bm{A}[\mathcal{J}],\bm{b}[\mathcal{J}]$ are the sub-matrix and sub-vector of $\bm{A},\bm{b}$.
@@
Accordingly, by substituting (21), (23), (24) into (20), the function between $\bm{z}_{d,\tau}^{*}$ at time-slot $\tau=2,...,T$ is,
@@
To sum up, the linear functions between the forecast $\hat{\bm{y}}_{d}$ and DA and RT primal solutions, which are defined in the neighborhood of $\hat{\bm{y}}_{d}$, are summarized in the (21)-(25). The function between the optimal dual solutions and the forecast $\hat{\bm{y}}_{d}$ is a constant function, whose output can be obtained by solving the dual problems of (10),(11), and (12). With these, we are ready to transform the upper-level objective (15a) to a function of the forecast $\hat{\bm{y}}_{d}$. The details are in the next subsection.
IV-B Loss Function Derivation
By substituting the functions (21),(23),(24) and dual solutions into the upper-level objective (15a), the loss function in the neighborhood of the forecast $\hat{\bm{y}}_{d}$ is,
$\ell_{d}(\hat{\bm{y}}_{d}):=-\bm{\sigma}_{d}^{*\top}(\bm{\psi}_{\text{DA},d}+\bm{F}_{\text{DA}}^{y}\hat{\bm{y}}_{d})-\bm{\nu}_{d,1}^{*\top}(\bm{\psi}_{\text{RT},d,1}+$
$\bm{F}_{\text{RT}}^{\prime p}f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})+\bm{F}_{\text{RT}}^{\prime+-}f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d}))$ (26a)
[…]
Since functions in (26b) are linear, and the dual decision policies are constant, the loss function $\ell_{d}(\hat{\bm{y}}_{d})$ in the neighborhood of the forecast $\hat{\bm{y}}_{d}$ is linear, and outputs the overall operating cost (4a) given the forecast $\hat{\bm{y}}_{d}$. Naturally, the derivative of $\ell_{d}(\hat{\bm{y}}_{d})$ w.r.t. the forecast $\hat{\bm{y}}_{d}$, i.e., $\frac{\partial\ell_{d}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, measures the marginal impact of the forecast on the overall cost and is a constant. It is,
where the derivatives $\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ in (27) are constants, which measure the marginal impact of the forecast on the primal solutions. We have the following proposition regarding the loss function.
Proposition 3.
Proof.
According to Theorem 1, (21),(23),(24) are piecewise linear functions defined across the entire space of $\hat{\bm{y}}_{d}$. Since the loss function is the linear transformation of (21),(23),(24), it is also a piecewise linear function. ∎
Since (21),(23),(24) are determined by different active constraint index sets $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau-1},\forall\tau=2,...,T$ of DA and RT market clearing, each piece of the loss function is related with different active constraint index sets we well. One way is to enumerate all possible active constraint index sets, and derive the corresponding piece of the loss function [23]. However, the enumeration can be computationally expensive, particularly when dealing with large-scale optimization.
Therefore, we propose to derive the pieces of the loss function in an on-demand manner. Concretely, during training, we only recalculate the piece of the loss function when encountering new active index sets. The details are given in the next Section.
V Solution Strategy
Figure 5: The structure of neural network.
VI Case Study
[…]
Figure 7: The illustration of IEEE 9-bus system.
Table I: Cost and technical data in DA and RT market of IEEE 9-Bus system.
$G_{1}$ $G_{2}$ $G_{3}$
Marginal generation cost $\bm{\rho}$ (\$/MW) 20 22 24
Minimum generation power $\underline{\bm{p}}$ (MW) 0 0 0
DA market Maximum generation power $\overline{\bm{p}}$ (MW) 150 200 270
Lower ramping limits $\underline{\bm{r}}$ (MW) -90 -80 -70
Upper ramping limits $\overline{\bm{r}}$ (MW) 90 80 70
Marginal cost $\bm{\rho}_{+}$ (\$/MW) 50 52 54
Marginal utility $\bm{\rho}_{-}$ (\$/MW) 18 16 14
Marginal opportunity loss for up-regulation $\bm{\rho}_{+}-\bm{\rho}$ (\$/MW) 30 30 30
RT Market Marginal opportunity loss for down-regulation $\bm{\rho}-\bm{\rho}_{-}$ (\$/MW) 2 6 10
Up-regulation limit $\overline{\bm{p}^{+}}$ (MW) 60 60 60
Down-regulation limit $\overline{\bm{p}^{-}}$ (MW) 60 60 60
We use a four-layer ResNet as the forecasting model, which has 256 hidden layer units. Its structure is described in Figure 5. The input context consists of the numeric weather prediction (i.e., the predicted wind speed and direction at 10m and 100m altitude) of each wind farm in the system. We use Root Mean Squared Error (RMSE) on the test set for assessing the forecast quality, and the average overall operating cost on the test set, as defined in (4a), for evaluating the operation value. Four benchmark models are used for comparison: Two quality-oriented (Qua-E and Qua-Q), one value-oriented point forecasting approach (OptNet), and a stochastic market clearing (Sto-OPT-P), i.e.,
  1. 1.
  2. 2.
  3. 3.
  4. 4.
VI-A The Operational Advantage
[…]
Table II: RMSE, average operating cost on the test set and the training/test time of the proposed approach and benchmarks.
Proposed Qua-E Qua-Q OptNet Sto-OPT-P
[…]
Training time per epoch 25 s 0.08 s - 136 s -
Test time 5 s 5 s 5 s 5 s 64min
VI-B The Sensitivity Analysis
VI-B1 Performance under Different Wind Power Penetration
[…]
VI-B2 Performance under Different Up-regulation Cost in RT
[…]
Table III: Average operating cost in different settings of up-regulation costs in RT. The first column gives the RT up-regulation costs and marginal opportunity loss in each setting. Inside the parenthesis, the up-regulation marginal costs are given on the left of the slash, and the marginal opportunity loss is given on the right.
Settings Proposed (\$) Qua-E (\$) Cost reduction (%)
High up- regulation marginal cost ($G_{1}$: 80/60, $G_{2}$: 82/60, $G_{3}$: 84/60) 85114 92486 8%
Low up- regulation marginal cost ($G_{1}$: 21/1, $G_{2}$: 23/1, $G_{3}$: 25/1) 81517 81677 0.2%
VI-C Extending to Larger Scale System
We test the proposed approach on the IEEE 39-bus system, with 21 loads, 2 wind farms at buses 38 and 39, and 10 generators. The illustration of the IEEE 39-bus system and parameters for generators are given in our Github repo [24]. The average operating costs of the proposed one and Qua-E are shown in Fig. 8(b). The proposed approach results in lower average operating costs than Qua-E. With increased wind power capacity in the IEEE 39-bus system, the advantage of the proposed approach in decreasing the overall costs is even more pronounced compared to the IEEE 9-bus system.
VII Conclusions
We propose a value-oriented renewable energy forecasting approach, for minimizing the expected overall operating cost in the existing deterministic market clearing framework. We analytically derive the loss function for value-oriented renewable energy forecasting in sequential market clearing. The loss function is proved to be piecewise linear when the market clearing is modeled by linear programs. Additionally, we provide the analytical gradient of the loss function with respect to the forecast, which leads to an efficient training strategy. In the case study, compared to quality-oriented forecasting approach trained by MSE, the proposed approach can reduce average operating cost on the test set to 2.9% for the IEEE 9-bus system. Such an advantage is more obvious under large wind power capacity and high up-regulation costs. Under high up-regulation marginal costs, our approach can reduce the cost by up to 8%. Future work will include derivation of the value-oriented loss function for market clearing modeled by other types of optimization programs, such as quadratic optimization and conic optimization.

### v3

@@ Appendix B Proof of Cost Recovery and Revenue Adequacy
In the following, we show how to derive the functions $f_{d,\tau}^{x}(\hat{\bm{y}}_{d})$, $f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})$, $f_{d,\tau}^{z}(\hat{\bm{y}}_{d})$, and $f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})$. We define an operator $\Pi_{\mathcal{J}}:h(\bm{x})\mapsto\tilde{h}(\bm{x})$, where $h(\bm{x})=\bm{A}\bm{x}+\bm{b}$ is a linear function with the parameter $\bm{A},\bm{b}$, and $\mathcal{J}$ is the row index subset of $\bm{A},\bm{b}$. The output $\tilde{h}(\bm{x})$ of the operator is also a linear function, where $\tilde{h}(\bm{x})=\bm{A}[\mathcal{J}]\bm{x}+\bm{b}[\mathcal{J}]$ and $\bm{A}[\mathcal{J}],\bm{b}[\mathcal{J}]$ are the sub-matrix and sub-vector of $\bm{A},\bm{b}$.
@@
Accordingly, by substituting (40), (42), (43) into (20), the function between $\bm{z}_{d,\tau}^{*}$ and forecast $\hat{\bm{y}}_{d}$ at $\tau=2,...,T$ is,
@@
To sum up, the linear functions between the forecast $\hat{\bm{y}}_{d}$ and DA and RT primal solutions, which are defined in the neighborhood of $\hat{\bm{y}}_{d}$, are summarized in the (40)-(44).
Figure 9: Average operating cost under different wind power capacities on the training set.
Figure 10: The training convergence curve.
Appendix D Results on the training set
To demonstrate our model is not overfitting, we also report the results on the training set, which can be found in Fig. 9. The improvement achieved by the proposed method over Qua-E on the training set is similar to that observed on the test set in Fig. 7. This demonstrates the model does not suffer overfitting. The results are reasonable since we train the forecasting model on a relatively large dataset with 7008 samples. Also, we use ResNet as the forecasting model, which avoids overfitting by using residual connections and allows the network to learn identity mappings and promote better gradient flow.
Additionally, to demonstrate the convergence of the forecasting model, we report the training convergence curve, which illustrates the change in the loss (21) over training iterations. Since (21) is a linear function composed of linear and constant parts, minimizing it is equivalent to minimizing the linear part. Therefore, the linear part itself can reflect the convergence process, which is shown in Fig. 10. The loss decreases and converges rapidly over iterations.